OpenStudy (anonymous):

what is the area of a right triangle with sides of 8ft, 15 ft, and 17ft?

OpenStudy (anonymous):

Do you know of @Hero 's formula

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Sorry. I just had to.

hero (hero):

I assure you, it's not my formula LOL

OpenStudy (anonymous):

So do you know the answer?

OpenStudy (solomonzelman):

A= (1/2)bh

hero (hero):

But before using Hero's formula, I'd make sure it's not a right triangle first.

OpenStudy (solomonzelman):

the legs are b and h

OpenStudy (isaiah.feynman):

A diagram of the triangle would be helpful!

OpenStudy (solomonzelman):

No it would not be

OpenStudy (solomonzelman):

The legs of the triangle are height and base. GO for it.

OpenStudy (anonymous):

It's a right triangle

hero (hero):

It would not be what @SolomonZelman ?

OpenStudy (anonymous):

Their is no diagram

OpenStudy (solomonzelman):

replying to Isaiah Feynaman

OpenStudy (solomonzelman):

Diagram is not needed.

hero (hero):

Performing \(a^2 + b^2 = c^2\) would confirm whether or not the given sides form a right triangle.

OpenStudy (anonymous):

It a right triangle -.-

OpenStudy (anonymous):

The question is : what is the area of a right triangle with sides of 8ft, 15 ft, and 17ft?

hero (hero):